OpenStudy (anonymous):

Differentiating and integration power series: find power series representation for g centered at 0 by differentiating or integrating the power series for f (perhaps more than once) give the interval of convergence for the resulting series. g(x)=ln(1-3x). Using g(x)=1/1-3x

OpenStudy (amistre64):

a power series is of the form \[\sum_{n=0}^{\inf}f^{(n)}(a)\frac{(x-a)^n}{n!}\]right?

OpenStudy (anonymous):

I don't know is this a formula.

OpenStudy (amistre64):

its a general setup yes; where a is the value we are centering around, and f^n(a) defines the value of the derivative at a

OpenStudy (anonymous):

Ok

OpenStudy (amistre64):

...value of the nth derivative at "a" ; that is

OpenStudy (anonymous):

So how can I solve this budd I am very confused

OpenStudy (amistre64):

well, since your not familiar with this notation, i have to know what methods youve tried to begin with. otherwise ill just confuse you more

OpenStudy (anonymous):

I think is1

OpenStudy (amistre64):

find power series representation for g centered at 0 by differentiating or integrating the power series for f there is nothing defined as an "f" in your posting give the interval of convergence for the resulting series. g(x)=ln(1-3x). Using g(x)=1/1-3x i know how to work an interval of convergence, but the information here is a little garbled.

OpenStudy (anonymous):

Well I saw a method in book but it was confusing as hell like it said you should take dx of 1+x^2+x^3+.... Then some how they come up with Series of k=0 as k is app 0 then (k+1)

OpenStudy (anonymous):

For f let me see one min

OpenStudy (amistre64):

sounds like im on the right track at least. should we construct the function of "f" from g(x) = ln(1-3x) by chance?

OpenStudy (anonymous):

Well according to theorem 9.5 it says let function of f be defined by the power series sum c_k(x-a)^k on its interval of convergence I.

OpenStudy (anonymous):

Then it says 1.) f is continues function on i

OpenStudy (anonymous):

Then 2.) the power series may be differentiated or integrated term by term, and the resulting power series converges to f prime (x) or integral of f(x)dx+c, respectively, at all points in the interior of I, where c is an arbitrary constant.

OpenStudy (amistre64):

to simplify matters, lets use ln(u) such that u=1-3x. when x=0, u=1 \[g=ln(u)\]\[g'=u^{-1}\]\[g''=-u^{-2}\]\[g'''=2u^{-3}\]\[g^{(4)}=-3*2u^{-4}\]\[g^{(5)}=4*3*2u^{-5}\]\[...\]\[g^{(n)}=(-1)^{(n+1)}(n-1)!*u^{-n}\]; and with u=1, that last u^(-n) part becomes trivial \[f(u)=\sum_{n=0}^{\infty}g^{(n)}(1)\frac{u^n}{n!}\] \[f(u)=\sum_{n=0}^{\infty}(-1)^{(n+1)}(n-1)!\frac{u^n}{n!}\] \[f(u)=\sum_{n=0}^{\infty}(-1)^{(n+1)}\frac{u^n}{n}\] its been ahile, so im not too sure how accurate my thought is, so let me check it out first

OpenStudy (anonymous):

Sure thanks for trying I will be waiting

OpenStudy (amistre64):

yeah, something went awry there. the set up for the series should be, and using the k notation to keep it consistent with your material: \[-\sum_{k=1}^{\infty}\frac{(-3)^k~(-x)^k}{k}\]

OpenStudy (anonymous):

So now I should just put the given ln(1-3x) in this formula

OpenStudy (amistre64):

no, this formula is the equivalent of ln(1-3x) between x=-1/3 and x=1/3

OpenStudy (anonymous):

So okay going back to what the question is asking what do u suggest me to do.

OpenStudy (anonymous):

I can't I got to sign up

OpenStudy (amistre64):

.... thats odd, wolframalpha is a free website.

OpenStudy (amistre64):

heres a screen shot; its good to know why this process is important. working with a poly is much simpler than the log.

OpenStudy (anonymous):

Ok saw that

OpenStudy (amistre64):

the setup for this would be the same as i did, but using the ln(1-3x) instead of the u-substitution :)

OpenStudy (anonymous):

I am sorry not to look for free ride but in general can you please tell me how to solve this step by step thanks

OpenStudy (amistre64):

you form a power series of the setup\[f(x)=g(0)+g'(0)x+g''(0)\frac{x^2}{2!}+g'''(0)\frac{x^3}{3!}+...+g^{(n)}(0)\frac{x^n}{n!}+...\] where you use the values of the successive derivatives \[g=ln(1-3x)~:~g(0)=0\]\[g'=-3*(1-3x)^{-1}~:~g'(0)=-3\]\[g''=-3*-1*-3*(1-3x)^{-2}~:~g''(0)=-9\]etc.... its a rather long and drawn out process

OpenStudy (amistre64):

once you have a general rule for how the function f(x) is being worked out, you can take the summation notation and take the limit of it as a ratio of n/(n-1) terms to find the interval of convergence

OpenStudy (amistre64):

... or at least that is how im reading the problem.

OpenStudy (anonymous):

Ok one min please

OpenStudy (anonymous):

Okay I got the terms now so what's next?

OpenStudy (amistre64):

write up a summation notation that defines the polynomial

OpenStudy (amistre64):

what terms did you come up with?

OpenStudy (anonymous):

The derivatives of the ln(1-3x) then entering the 0 in them then getting ans.

OpenStudy (amistre64):

do they match with these? 0, -3, -9, -54, -486, ...

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

good, then at least ive got that correctly :)

OpenStudy (amistre64):

soo; for the g' s the rule:\[(-1)^{k+1}~(-3)^k~(k-1)!\]and we apply this to the \[\frac{x^k}{k!}\]to get: \[\sum(-1)^{k+1}~(-3)^k~(k-1)!\frac{x^k}{k!}\] \[\sum(-1)^{k+1}~(-3)^k~(k-1)!\frac{x^k}{k!}\] \[\sum(-1)^{k+1}~(-3)^k~\frac{x^k}{k}\]and to clean it up \[\sum~\frac{(-1)^{k+1}~(-3)^k~x^k}{k}\]is what we get to work with to find the interval of convergence

OpenStudy (amistre64):

we find the convergence by taking the limit of the ratio of k/k-1 as k approaches infinity \[lim ~\frac{(-1)^{k+1}~(-3)^k~x^k}{k}\frac{k-1}{(-1)^{k}~(-3)^{k-1}~x^{k-1}}\] \[lim ~\frac{(-1)~(-3)~x}{k}\frac{k-1}{1}\] \[lim ~(-1)~(-3)~x\frac{k-1}{k}\] since everything without a k in it after the simplification is over has no effect on the limit; take it to the outside and wrap it in absolute values \[ |(-1)(-3)x|\lim_{k\to\ inf}|\frac{k-1}{k}\] \[ |3x|\lim_{k\to\ inf}|\frac{k-1}{k}\] the limit as kapproaches infinity of (k-1)/k is 1, leaving us with |3x|

OpenStudy (amistre64):

now for reasons I cant recall, this is defined as the radius of convergence; and the interval takes this |3x| and compares it against 1 for the interval of convergence: \[|3x|<1\]\[3|x|<1\]\[|x|<\frac13\]or simply giving as interval of convergence as\[-\frac13<x<\frac13\]

OpenStudy (anonymous):

allow me to butt in i think you were supposed to integrate the series for \(\frac{1}{1-3x}\) term by term. of course computing directly will work as well

OpenStudy (anonymous):

Satellite 73 can you please explain further.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

start with the well known fact \[\frac{1}{1-x}=1+x+x^2+x^3+...=\sum_{k=1}^{\infty}x^k\] for \(|x|<1\)

OpenStudy (anonymous):

this is the standard trick, and if you are confused at that step (it looks like you were above) then you have to take that as a given but since you have \(\frac{1}{1-3x}\) you have to replace \(x\) by \(3x\) giving [\frac{1}{1-3x}=1+3x+9x^2+...=\sum_{k=0}^{\infty}(3x)^k\]

OpenStudy (anonymous):

hmm \[\frac{1}{1-3x}=1+3x+9x^2+...=\sum_{k=0}^{\infty}(3x)^k\] thats better

OpenStudy (anonymous):

then you integrate this sucker term by term, since it is almost the derivative of \[\ln(1-3x)\]

OpenStudy (anonymous):

How for the last step?

OpenStudy (anonymous):

after you integrate term by term (i.e. write a formula for what you would get) multiply all by \(-3\) because \[\frac{d}{dx}\ln(1-3x)=\frac{-3}{1-3x}\] by the chain rule

OpenStudy (anonymous):

I know budd but how can integrate this

OpenStudy (anonymous):

it is just another way of doing it, using the gimmick of knowing the power series of the derivative, (you have to know this for the trick to work) and then integrating term by term

OpenStudy (anonymous):

\[1\to x\] \[3x\to \frac{3x^2}{2}\] \[(3x)^2\to \frac{3^2x^3}{3}\] \[(3x)^3\to \frac{3^3x^4}{4}\] literally term by term it really amounts to figuring out the general term that is all

OpenStudy (anonymous):

Oh I got you budd you wanted me to put the values back in the integration formula so now what budd

OpenStudy (anonymous):

multiply by -3 and you are done

OpenStudy (anonymous):

So i got -3x-9x^2/3-9x^3

OpenStudy (anonymous):

Oh also what about interval of convergence?